OpenStudy (anonymous):

A hardware store has the policy of marking up their goods by 85%. Calculate the cost price if the item is sold for $3.70

OpenStudy (anonymous):

is it 3.70/1.85 =2

OpenStudy (anonymous):

or better 3.70*100/(100+85)=2

OpenStudy (anonymous):

$3.70 / (1+0.85) = $2 CHECK: $2.00 * 1.85 = $3.70